A person walks 24 m East and then walks
40 m at an angle 28◦ North of East. What is the magnitude of the total displacement? Answer in units of m.
1
Expert's answer
2010-09-29T11:56:55-0400
We have two displacements: r1, r2 ant total displacement r.
r1x=24m
r2x=40×cos(28°)=35.3179 m
rx=r1x+r2x=24+35.3179=59.3179 m

r1y=0
r2y=40×sin(28°)=18.7789 m
ry=r1y+r2y=0+18.7789=18.7789 m

Hence,
r=√rx²+ry²=√59.3179²+18.7789²=62.2185 m